# Portable, Massively Parallel Implementation of a Material Point Method for Compressible Flows

### v2

@@ 2.2 Numerical Method
The link between the Eulerian grid and the Lagrangian particles is given by the Particle-to-Grid (P2G) and Grid-to-Particle (G2P) mappings, which are the building blocks of the computational algorithm, and that can be derived during the discretisation of the weak form (3), that leads to a semi-discrete equation of the form [^1] Here and in the following, when it is useful to distinguish, grid quantities are written in lower case Roman characters, while particle ones are written in either upper case Roman or Greek characters; moreover, $i$ denotes grid nodes index, while $k$ denotes particles index.: $\forall i=0\ldots N_{h}-1\quad m_{i}\mathbf{\dot{v}}_{i}=\mathbf{f}_{i}$ (6) where $N_{h}$ denotes the number of grid nodes.
The latter can be obtained following the standard MPM procedure (see e.g. [51]), including five main steps: 1. discretise the material domain into a mesh of material sub-domains, with mass concentrated at material points $\rho(\mathbf{x},t)=\sum_{k=0}^{N_{p}-1}M_{k}\delta(\mathbf{x}-\mathbf{X}_{k})$ (7) where $N_{p}$ denotes the total number of material points, and insert (7) into (3); 2. discretise space with a standard Finite Elements Method (FEM) grid; 3. approximate any field and test function relying on FE grid shape functions $u_{i}$; 4. apply arbitrariness of test function to get the semi-discrete equation; 5. possibly, lump the mass matrix $m_{i}\!=\!\sum_{j}m_{ij}\!=\!\sum_{j}\sum_{k}M_{k}u_{i}(\mathbf{X}_{k})u_{j}(\mathbf{X}_{k})\!=\!\sum_{k}M_{k}u_{i}(\mathbf{X}_{k})$ (8) where $u_{i}(\mathbf{X}_{k})$ denotes the basis function of the solution space and test functions space centred on the $i$-th grid node and evaluated in the $k$-th particle position, $\mathbf{X}_{k}=(X_{k},Y_{k})$.
It is worth noting that it is possible to derive the same formulation also by adopting the standard Galerkin-Finite Element procedure, provided material points are considered as moving quadrature points [19]. In fact, standard MPM can be seen as a modification of traditional FEM, in which the integration points are not related to the Finite Element mesh and are Lagrangian, and thus suitable for history dependent models too [14].
Following this method, we start from the motion equation expressed in weak formulation, which reads “find $\mathbf{v}\in V$ such that equation (3) holds $\forall\mathbf{w}\in V$”, $V$ being a proper Hilbert space. To discretise the problem, we introduce a finite dimensional sub-space $V_{h}=span\{u_{i}(x,y)\}$ , project equation (3) onto $V_{h}$, and, since it has to hold for any $\mathbf{w}_{h}\in V_{h}$, we find $\mathbf{v}_{h}$ such that equation (3) holds for each function in $V_{h}$ base.
Since $V_{h}$ is finite dimensional, it is possible to express any function in terms of finite sum, e.g.
$v_{h}(x,y)=\sum_{i}v_{i}u_{i}(x,y)$ (9)
In other words, we write equation (3) choosing as test function the base functions $u_{i}$ centred on the $i-th$ grid node, and we solve numerically the integrals by means of quadrature rules, evaluating the integral at a finite set of points, called the integration points $p$, and perform the weighted sum
@@
If the material points are taken as integration points, that is, if basis functions are evaluated in the $k-th$ particle position, with weights being their volumes (i.e. $\omega_{p}=\tfrac{M_{k}}{\rho_{k}}$), then FEM equations reduces to MPM ones. In particular,
$\mathbf{f}^{internal}_{i}=-\int_{\Omega}\bm{\sigma}\nabla u_{i}d\Omega\simeq-\sum_{k=0}^{N_{p}-1}\bm{\sigma}_{k}\tfrac{M_{k}}{\rho_{k}}\nabla u_{i}(X_{k},Y_{k})$ (11)
which is the internal component of the force in the semi-discrete equation (6), where we have already applied mass matrix lumping.
In particular, equation (11) expresses the Particle-to-Grid mapping of internal forces, which, in our inviscid gas model with $\bm{\sigma}=-p\mathbb{I}$, maps the pressure force from the material points to the grid nodes. To distinguish this from the Particle-to-Grid which does not involve moving the derivative to the test function, we call operations as (11) P2GD.
P2GD is one of the four building blocks of MPM algorithm, the others being:
  • P2G, without weak derivatives, which is needed since grid variables are reset at every time step (see sec. 3);
  • G2P, which maps quantities back to the particles after equations have been solved numerically;
  • G2PD, which also computes in-place a weak derivative.
In this work, we define these operations, which we call kernels following GPU literature, in a general way to underline their computational structure for later implementation and performance analysis, even though they are applied only as written in the algorithm in section 3. More details on how these operations are derived can be found in the original papers for the standard MPM [51], in [64, 65] for a first application to compressible fluids, and in [14, 63, 19] for later reviews.
Denote by $\mathbf{a}=[a_{i}]\in\mathbb{R}^{N_{h}}$ the array of values of some quantity $a(x,y)$ at the grid vertices $(x_{i},y_{i}),i=0\ldots N_{h}-1$, and let $\mathbf{A}=[A_{k}]\in\mathbb{R}^{N_{p}}$ be the array of values of some quantity $A(X,Y)$ at the material points $(X_{k},Y_{k}),k=0\ldots N_{p}-1$.
Particle-to-Grid and Grid-to-Particle mappings read:
$a_{i}=\mathrm{P2G}(\mathbf{A};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}A_{k}\ u_{i}(X_{k},Y_{k}),\quad i=0\ldots N_{h}-1$ (12)
$A_{k}=\mathrm{G2P}(\mathbf{a};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}a_{i}\ u_{i}(X_{k},Y_{k}),\quad k=0\ldots N_{p}-1$ (13)
$\left(\nabla a\right)_{i}=\mathrm{P2GD}(\mathbf{A};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}A_{k}\ \left(\nabla u\right)_{i}(X_{k},Y_{k}),\quad i=0\ldots N_{h}-1$ (14)
$\left(\nabla A\right)_{k}=\mathrm{G2PD}(\mathbf{a};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}a_{i}\ \left(\nabla u\right)_{i}(X_{k},Y_{K}),\quad k=0\ldots N_{p}-1.$ (15)
@@
It is useful to generalise these kernels to the case where the mapped quantity is not stored in memory but computed on the fly as a function of other state variables, as in equation (11). To this end, let $\mathcal{F}:\mathbb{R}^{m}\to\mathbb{R}$, and rewrite the above kernels as follows :
$a_{i}=\mathrm{P2G}(\mathcal{F},\mathbf{A}^{(0)},\ldots,\mathbf{A}^{(m-1)};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}\mathcal{F}\left({A}_{k}^{(0)},\ldots{A}_{k}^{(m-1)}\right)\ u_{i}(X_{k},Y_{k}),\quad i=0\ldots N_{h}-1$ (15)
$A_{k}=\mathrm{G2P}(\mathcal{F},\mathbf{a}^{(0)},\ldots,\mathbf{a}^{(m-1)};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}\mathcal{F}\left({a}_{i}^{(0)},\ldots{a}_{i}^{(m-1)}\right)\ u_{i}(X_{k},Y_{k}),\;k=0\ldots N_{p}-1$ (16)
@@
$\left(\nabla a\right)_{i}=\mathrm{P2GD}(\mathcal{F},\mathbf{A}^{(0)},\ldots,\mathbf{A}^{(m-1)};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}\mathcal{F}\left({A}_{k}^{(0)},\ldots,{A}_{k}^{(m-1)}\right)\ \left(\nabla u\right)_{i}(X_{k},Y_{k}),\\ i=0\ldots N_{h}-1$ (17)
$\left(\nabla A\right)_{k}=\mathrm{G2PD}(\mathcal{F},\mathbf{a}^{(0)},\ldots,\mathbf{a}^{(m-1)};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}\mathcal{F}\left({a}_{i}^{(0)},\ldots{a}_{i}^{(m-1)}\right)\ \left(\nabla u\right)_{i}(X_{k},Y_{K}),\\ k=0\ldots N_{p}-1.$ (18)
@@
For the numerical discretisation of the Eulerian aspects of the method, we choose a Cartesian grid of quadrilateral elements. Among the possible choices, this one lends itself naturally to computationally efficient refinements through a quad-tree approach [46]. Moreover, we opt for Lagrangian hat functions as basis for the functional spaces of both the solution and the test function, setting our treatment in a $\mathbb{Q}_{1}$ framework. [^2] In sec. 5 we discuss the extension to higher order basis functions.
@@
This approach corresponds to bi-linear weighting function of standard MPM, and is known to cause cell-crossing error and spurious oscillations in the computed quantities. The reason can be easily understood by the above equations: when $u$ is (bi)-linear, $\nabla u$ is constant, but potentially different in neighbouring cells. Thus, when crossing cells, particles might experience discontinuous oscillations. This issue can be numerically tackled with many strategies, among which adding artificial viscosity [63] or choosing higher order basis functions. For simplicity of implementation, we choose the first, leaving the second for a future work.
@@
With respect to standard MPM [64, 14], the present algorithm adds a step to take into account the flow around obstacles. Namely, we subtract the inward component of both linear momentum and force from the grid nodes whose cell is marked as an obstacle cell in following equations (stemming from the method for granular materials [3]):
$(m\mathbf{v}_{i}^{o})^{\prime}=m\mathbf{v}_{i}^{o}-\dfrac{m\mathbf{v}_{i}^{o}\cdot\mathbf{n}_{i}^{o}-|m\mathbf{v}_{i}^{o}\cdot\mathbf{n}_{i}^{o}|}{2}\mathbf{n}_{i}^{o}$ (19)
$(\mathbf{f}_{i}^{o})^{\prime}=\mathbf{f}_{i}^{o}-\dfrac{\mathbf{f}_{i}^{o}\cdot\mathbf{n}_{i}^{o}-|\mathbf{f}_{i}^{o}\cdot\mathbf{n}_{i}^{o}|}{2}\mathbf{n}_{i}^{o}$ (20)
@@
where the subscript denotes the $i-th$ grid node, the superscript $o$ stands for obstacle and $\mathbf{n}$ is the normal unit vector, positive if directed outward from the obstacle.
@@
$sd(\mathbf{x})=\begin{cases}-\inf_{y\in\partial\Omega}d(\mathbf{x},\mathbf{y})\quad\mathbf{x}\in\Omega\\ \inf_{y\in\partial\Omega}d(\mathbf{x},\mathbf{y})\quad\mathbf{x}\notin\Omega\end{cases}$ (21)
where $d(x,y)$ is the Euclidean distance. Taking the gradient yields [^3] An algorithm for signed distance function computation can be found in [10], and an example in a machine learning based framework that implicitly represents geometries from noisy surface points can be found in [13].
@@
where $\alpha$ collects constants. The range of admissible time steps is bounded from above by the ratio between the smallest grid cell dimension $\Delta x_{min}$ and the highest speed $v_{max}$, considering both particles velocity and speed of sound $c_{s}=\sqrt{\gamma\tfrac{p}{\rho}}$. The efficiency cost of enforcing small time steps is balanced by the need of resolving fast dynamics. In addition, explicit-in-time methods sidestep the need to solve global linear systems at each time step, thus enjoying better scalability properties than implicit methods.
@@ 3 Portable MPM Implementation
This section describes the algorithm that was presented in section 2 and how it was implemented to ensure portability on different architectures and efficiency on GPUs. The development of the code benefited in part from publicly available Open Source libraries: Thrust [^4] https://github.com/NVIDIA/cccl/tree/main/thrust, JSON [^5] https://github.com/nlohmann/json, and quadgrid [^6] https://github.com/carlodefalco/quadgrid. In detail, the algorithm implementation is composed of the following steps:
@@
    Data generation: the creation of simulations input is performed separately from code execution. For simple geometries it is handled via ad-hoc C++ scripts; for complex geometries, requiring a careful computation of the signed distance function, an Octave script is used instead, to take advantage of Octave built-in functions. In both cases, since objects are considered at rest, normal unit vectors are computed only once, and boundary cells are marked with a flag corresponding to the boundary condition. Finally, the initial data file is saved in JSON format.
@@
    The time advancing loop starts; inner steps, except from output operation, are implemented as Thrust or STL for _eachs, depending on the architecture chosen at compile time.
@@
      Grid variables are reset and current grid variables are overwritten with the ones saved during construction stage, performing a device to device copy. This step can be performed since all history dependence is maintained by Lagrangian particles.
@@
      Equations (12) and (17) are applied to map particles mass, momentum and force to the grid; during force P2GD, artificial viscosity is applied as in (5).
@@
      Grid nodes momentum is advanced in time.
@@
      Velocity and acceleration generalised G2P (16) is applied, and next, velocity is updated. Acceleration itself is not stored, but computed as the result coming from a local G2P of the relevant nodes ratio between force and mass; moreover, G2P kernels are performed on particles index, as shown in the snippet in Fig. 1, since this has proved to provide some acceleration with respect to an approach based on grid nodes index. Following standard MPM [14], we store the G2P velocity in a temporary velocity vector $\mathbf{\overline{V}}$ and don’t use it to update particles velocity; however, we postpone its usage.
@@
      Particles strain rate tensor (velocity gradient) is computed as the G2PD (18) of the ratio between grid nodes momentum and mass, and then particles properties are updated; also in this case, the particle index is used for the G2PD kernel, and artificial viscosity is applied in the update of internal energy density.
@@
      Particles position are updated using $\mathbf{\overline{V}}$. Thus, the time advancing scheme is the same of [14]. However, particles moving is postponed, so that, in all the above steps, basis functions and their gradient can be computed on the fly, without the need to store the value they had at former particles position.
@@
Algorithm 1 summarises the for loop procedure.
@@
$\qquad m=\mathrm{P2G}(M),\ m\mathbf{v}=\mathrm{P2G}(M\mathbf{V}),\ \mathbf{f}=\mathrm{P2GD}((P+Q)\tfrac{M}{\rho})$
4:     Boundary conditions enforcement on $m\mathbf{v},\,\mathbf{f}$ obstacle treatment as in (19), (20)
@@
$\qquad(m\mathbf{v})^{\prime}=m\mathbf{v}+\Delta t\mathbf{f}$
@@
$\qquad\overline{\mathbf{V}}=\mathrm{G2P}(\tfrac{m\mathbf{v}}{m}),\ {\mathbf{A}}=\mathrm{G2P}(\tfrac{\mathbf{f}}{m})$
@@
$\qquad\mathbf{V}^{\prime}=\mathbf{V}+\Delta t\mathbf{A}$

### v4

@@ 2.2 Numerical Method
In this subsection we derive the formulae which serve as Particle-to-Grid mappings from the discretisation of the weak form of the motion equation (8). Then, building upon the derived expression, we define four mapping operations in a general way, such that they can be applied to different quantities to formulate the algorithm.
This step is indeed fundamental, since by expressing the whole algorithm in terms of four main kernels we are then able to perform better computational analysis, measures and optimisation.
In the following, we proceed in the afore-mentioned Galerkin way, as discussed also in [13] and done by [52, 53].
When it is useful to distinguish, grid quantities are written in lower case Roman characters, while particle quantities are written in either upper case Roman or Greek characters. Moreover, $i$ denotes grid nodes index, while $k$ denotes particles index.
To discretise (8), we introduce a finite dimensional sub-space of $V$, which we call $V_{h}$, and a basis $\{\phi_{i}(x,y)\}_{i=0}^{N_{h}-1}$, such that $V\ni V_{h}=span\{\phi_{i}(x,y)\}_{i=0}^{N_{h}-1}$. Then, we project equation (8) onto $V_{h}$, and, since $V_{h}$ is finite dimensional, it is possible to express any function in terms of finite sum, e.g.
$\mathbf{\dot{v}}_{h}(x,y)=\sum_{j=0}^{N_{h}-1}\mathbf{\dot{v}}_{j}\phi_{j}(x,y).$ (9)
Moreover, since the semi-discrete (8) has to hold for any $\mathbf{w}_{h}\in V_{h}$, we can choose as test function any basis function. We thus get the semi-discrete weak formulation [^1] By taking both $\mathbf{v_{h}},\mathbf{w_{h}}\in V_{h}\subset[H_{0}^{1}(\Omega)]^{2}$, as customary in Galerkin method, we are implicitly assuming homogeneous Dirichlet boundary conditions. When this is not the case, it is enough to introduce a lifting operator, see e.g. [17].:
Find $\mathbf{v_{h}}\in V_{h}$ such that $\forall i=0:N_{h}-1$
@@
Integrating numerically (10) with a quadrature rule having particles volumes as weights and particles position as integration points, we can express it in algebraic form:
$\forall i=0\ldots N_{h}-1\quad\sum_{j=0}^{N_{h}-1}m_{ij}\mathbf{\dot{v}}_{j}=\mathbf{f_{i}^{\,int}}$ (11)
having introduced the mass matrix $[m_{ij}]$ s.t.
$m_{ij}=\sum_{k=0}^{N_{p}-1}M_{k}\phi_{i}(X_{k},Y_{k})\phi_{j}(X_{k},Y_{k})$ (12)
and the internal force vector
$\mathbf{f}_{i}=\sum_{k=0}^{N_{p}-1}(P_{k}+Q_{k})\frac{M_{k}}{\rho_{k}}\nabla\phi_{i}(X_{k},Y_{k})$ (13)
where $k$ is the particles index, $N_{p}$ the particles number, $M_{k}$ is the $k-th$ particle mass such that its density is $\rho_{k}=\tfrac{M_{k}}{V_{k}}$, $V_{k}$ being its volume used as quadrature weight, $P_{k}$ and $Q_{k}$ are the $k-th$ particle pressure and artificial pressure, $\phi_{i}(X_{k},Y_{k})$ the $V_{h}$ base function centred on the $i-th$ grid node and evaluated at the $k-th$ particle position and $\nabla\phi_{i}(X_{k},Y_{k})$ the evaluation at $(X_{k},Y_{k})$ of the gradient of $\phi_{i}(x,y)$.
In order to avoid to solve a linear system, as customary in MPM [13, 52], we lump the mass matrix and define the nodal mass matrix whose diagonal terms are
$m_{i}=\sum_{k=0}^{Np-1}M_{k}\phi_{i}(X_{k},Y_{k})$ (14)
We thus have obtained an algebraic system in the form
$\forall i=0\ldots N_{h}-1\quad m_{i}\mathbf{\dot{v}}_{i}=\mathbf{f_{i}^{\,int}}$ (15)
where $[m_{i}]$ is the matrix of Eulerian grid nodal mass, which corresponds to the mass-lumped version of the original MPM algorithm mass matrix [57, 13, 53], and $\mathbf{\dot{v}}_{i}$ is the unknown vector of nodal acceleration. More precisely, the expressions just obtained for $m_{i}$ and $\mathbf{f_{i}^{\,int}}$ are respectively equations (3.1) and (3.3) in [13], assuming that our model (5a) is taken into account.
The so defined $m_{i}$ and $\mathbf{f_{i}^{\,int}}$ expressions constitute examples of what we call P2G and P2GD kernels, i.e. mapping operations from the Lagrangian particles to the Eulerian grid, where we the D stands for the derivative operator applied to the basis function, coming from having moved it from the stress to the test function in the weak formulation.
As anticipated, to better analyse, measure and optimise the computational performance of the algorithm, it is important to define such mapping operations in general terms, to be able both to stress their computational features and to apply the very same kernels to different quantities as specified in the algorithm of [13].
@@
In order to do so, denote by $\mathbf{a}=[a_{i}]\in\mathbb{R}^{N_{h}}$ the array of values of some quantity $a(x,y)$ at the grid vertices $(x_{i},y_{i}),i=0\ldots N_{h}-1$, and let $\mathbf{A}=[A_{k}]\in\mathbb{R}^{N_{p}}$ be the array of values of some quantity $A(X,Y)$ at the material points $(X_{k},Y_{k}),k=0\ldots N_{p}-1$. It is useful to generalise the kernels to the case where the mapped quantity is not stored in memory but computed on the fly as a function of other state variables, as in equation (11).
To do so, let $\mathcal{F}:\mathbb{R}^{m}\to\mathbb{R}$ and $\mathcal{G}:\mathbb{R}^{m}\to\mathbb{R}^{2}$; we define Particle-to-Grid and Grid-to-Particle kernels as follows:
$\mathrm{P2G}(\mathcal{F},\mathbf{A}^{(0)},\ldots,\mathbf{A}^{(m-1)};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}\mathcal{F}\left({A}_{k}^{(0)},\ldots{A}_{k}^{(m-1)}\right)\ \phi_{i}(X_{k},Y_{k}),\;i=0\ldots N_{h}-1$ (16)
$\mathrm{G2P}(\mathcal{F},\mathbf{a}^{(0)},\ldots,\mathbf{a}^{(m-1)};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}\mathcal{F}\left({a}_{i}^{(0)},\ldots{a}_{i}^{(m-1)}\right)\ \phi_{i}(X_{k},Y_{k}),\;k=0\ldots N_{p}-1$ (17)
@@
$\mathrm{P2GD}(\mathcal{G},\mathbf{A}^{(0)},\ldots,\mathbf{A}^{(m-1)};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}\mathcal{G}\left({A}_{k}^{(0)},\ldots,{A}_{k}^{(m-1)}\right)\cdot\nabla\phi_{i}(X_{k},Y_{k}),\;i=0\ldots N_{h}-1$ (18)
$\mathrm{G2PD}(\mathcal{F},\mathbf{a}^{(0)},\ldots,\mathbf{a}^{(m-1)};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}\mathcal{F}\left({a}_{i}^{(0)},\ldots{a}_{i}^{(m-1)}\right)\nabla\phi_{i}(X_{k},Y_{K}),\;k=0\ldots N_{p}-1.$ (19)
With these definitions, the motion equation can be solved on the Eulerian grid starting from Lagrangian particles quantities, assuming that proper boundary and initial condition are applied. Indeed, equation (15) can be formulated employing the above defined general kernels (16), (18) as:
$\forall i=0\ldots N_{h}-1,$
[…]
where:
$\mathbb{I}:\mathbb{R}\rightarrow\mathbb{R},\,\mathbb{I}(x)=x$
$\zeta:\mathbb{R}^{3}\times\mathbb{R}_{0}\rightarrow\mathbb{R}^{2},\,\zeta(x,y,z,t)=\begin{pmatrix}x+y\\ 0\end{pmatrix}\frac{z}{t}$
$\eta:\mathbb{R}^{3}\times\mathbb{R}_{0}\rightarrow\mathbb{R}^{2},\,\eta(x,y,z,t)=\begin{pmatrix}0\\ x+y\end{pmatrix}\frac{z}{t}$
In the next section the detailed implementation of the algorithm is illustrated, together with the expression of computational steps according to the kernels formalism, that allow to obtain the original MPM equation as in [13]
@@
For the numerical discretisation of the Eulerian aspects of the method, we choose a Cartesian tensor product grid of quadrilateral elements. Among the possible choices, this one lends itself naturally to computationally efficient refinements through a quad-tree approach [49, 13]. Moreover, we opt for $\mathbb{Q}_{1}$ FEM basis functions for the functional spaces of both the solution and the test function, see e.g. [17]. [^2] In section 5 we discuss the extension to higher order basis functions.
@@
This approach corresponds to bi-linear weighting function of standard MPM, and is known to cause cell-crossing error and spurious oscillations in the computed quantities. The reason can be easily understood by the above equations: when $\phi$ is (bi)-linear, $\nabla\phi$ is constant, but potentially different in neighbouring cells. Thus, when crossing cells, particles might experience discontinuous oscillations. This issue can be numerically tackled with many strategies, among which adding artificial viscosity [68] or choosing higher order basis functions. For simplicity of implementation, we choose the first, leaving the second for a future work.
@@
With respect to standard MPM [69, 13], the present algorithm adds a step to take into account the flow around obstacles. Namely, we subtract the inward component of both linear momentum and force from the grid nodes whose cell is marked as an obstacle cell in following equations (stemming from the method for granular materials [3]).
$\forall i=0\ldots N_{h}-1$:
${(\mathbf{mv})_{i}^{o}}^{\prime}=(\mathbf{mv})_{i}^{o}-\dfrac{(\mathbf{mv})_{i}^{o}\cdot\mathbf{n}_{i}^{o}-|(\mathbf{mv})_{i}^{o}\cdot\mathbf{n}_{i}^{o}|}{2}\mathbf{n}_{i}^{o}$ (21)
${\mathbf{f}_{i}^{o}}^{\prime}=\mathbf{f}_{i}^{o}-\dfrac{\mathbf{f}_{i}^{o}\cdot\mathbf{n}_{i}^{o}-|\mathbf{f}_{i}^{o}\cdot\mathbf{n}_{i}^{o}|}{2}\mathbf{n}_{i}^{o}$ (22)
@@
where $\mathbf{mv}$ is the linear momentum, $\mathbf{f}$ the internal force, the subscript denotes the $i-th$ grid node, the superscript $o$ stands for obstacle and $\mathbf{n}$ is the normal unit vector, positive if directed outward from the obstacle.
@@
$sd(\mathbf{x})=\begin{cases}-\inf_{\mathbf{y}\in\partial\Omega}d(\mathbf{x},\mathbf{y})\quad\mathbf{x}\in\Omega\\ \inf_{\mathbf{y}\in\partial\Omega}d(\mathbf{x},\mathbf{y})\quad\mathbf{x}\notin\Omega\end{cases}$ (23)
where $d(\mathbf{x},\mathbf{y})$ is the Euclidean distance. Taking the gradient yields [^3] An algorithm for signed distance function computation can be found in [9], and an example in a machine learning based framework that implicitly represents geometries from noisy surface points can be found in [12].
@@
where $\alpha$ collects constants. The range of admissible time steps is bounded from above by the ratio between the smallest grid cell dimension $\Delta x_{min}$ and the highest speed $v_{max}$, considering both particles velocity and speed of sound, given by the square root of the adiabatic index $\gamma$ multiplied by the ratio of gas pressure and mass density. The efficiency cost of enforcing small time steps is balanced by the need of resolving fast dynamics. In addition, explicit-in-time methods sidestep the need to solve global linear systems at each time step, thus enjoying better scalability properties than implicit methods.
@@ 3 Portable MPM Implementation
This section describes how we formulated the MPM algorithm of [13] within the framework that was presented in section 2 and how we implemented it to ensure portability on different architectures and efficiency on GPUs. The development of the code benefited in part from publicly available Open Source libraries: Thrust [^4] https://github.com/NVIDIA/cccl/tree/main/thrust, rocThrust [^5] https://github.com/ROCm/rocThrust, “JSON for Modern C++” [^6] https://github.com/nlohmann/json, and quadgrid. [^7] https://github.com/carlodefalco/quadgrid
Let $\mathbb{I},\zeta,\eta$ denote the fields defined in 2 and let $\theta$ be:
$\theta:\mathbb{R}\times\mathbb{R}_{0}\rightarrow\mathbb{R},\,\theta(x,y)=\frac{x}{y}$
In detail, the algorithm implementation is composed of the following steps:
@@
    Data generation: the creation of simulations input is performed separately from code execution. For simple geometries it is handled via ad-hoc C++ scripts; for complex geometries, requiring a careful computation of the signed distance function, an Octave [^8] https://octave.org/ script is used instead, to take advantage of Octave built-in functions. In both cases, since objects are considered at rest, normal unit vectors are computed only once, and boundary cells are marked with a flag corresponding to the boundary condition. Finally, the initial data file is saved in JSON format.
@@
    The time advancing loop starts; inner steps, except from output operation, are implemented as Thrust or STL for_eaches, depending on the choice made at compile time.
@@
      Grid variables are reset and current grid variables are overwritten with the ones saved during construction stage, performing a device to device copy. This step can be performed since all history dependence is maintained by Lagrangian particles. Specifically, the variables that have been reset to their original input values are vectors holding grid nodal mass, $x$ and $y$ components of linear momentum, force normal to the obstacle’s unit vector, as well as markers of whether such node has to be considered on a domain boundary or inside an obstacle.
@@
      Equations (16) and (18) are applied to map particles mass, momentum and force to the grid:
      $\forall i=0,\ldots,N_{h}-1$
      $m_{i}=\mathrm{P2G}(\mathbb{I},\mathbf{M};x_{i},y_{i})$
      $(mv_{x})_{i}=\mathrm{P2G}(\mathbb{I},\mathbf{MV_{x}};x_{i},y_{i})$
      $(mv_{y})_{i}=\mathrm{P2G}(\mathbb{I},\mathbf{MV_{y}};x_{i},y_{i})$
      $(f_{x})_{i}=\mathrm{P2GD}(\zeta,{\mathbf{P,Q,M,\bm{\rho}}};x_{i},y_{i})$
      $(f_{y})_{i}=\mathrm{P2GD}(\eta,{\mathbf{P,Q,M,\bm{\rho}}};x_{i},y_{i})$
@@
      Grid nodes momentum is advanced in time. [^9] In the implementation we use the linear momentum, as in eq. (3.5) of [13], instead of the equivalent formulation in terms of mass and velocity as in equation (2.2) or as in eq. (3.23) of [52].
      $\forall i=0,\ldots N_{h}-1$
      $(mv_{x})_{i}^{\prime}=(mv_{x})_{i}+\Delta t(f_{x})_{i}$
      $(mv_{y})_{i}^{\prime}=(mv_{y})_{i}+\Delta t(f_{y})_{i}$
@@
      Velocity and acceleration generalised G2P (17) is applied, and next, velocity is updated. Acceleration itself is not stored, but computed as the result coming from a local G2P of the relevant nodes ratio between force and mass. Following standard MPM [13], we store the G2P velocity in a temporary velocity vector $\mathbf{\overline{V}}$ and don’t use it to update particles velocity; however, we postpone its usage.
      $\forall k=0,\ldots N_{p}-1$
      $(\overline{V}_{x})_{k}=\mathrm{G2P}(\theta,\mathbf{mv_{x}},\mathbf{m};X_{k},Y_{k})$
      $(\overline{V}_{y})_{k}=\mathrm{G2P}(\theta,\mathbf{mv_{y}},\mathbf{m};X_{k},Y_{k})$
      $(A_{x})_{k}=\mathrm{G2P}(\theta,\mathbf{f_{x}},\mathbf{m};X_{k},Y_{k})$
      $(A_{y})_{k}=\mathrm{G2P}(\theta,\mathbf{f_{y}},\mathbf{m};X_{k},Y_{k})$
@@
      $\forall i=0,\ldots N_{h}-1$
      $(mv_{x})_{i}=\mathrm{P2G}(\mathbb{I},\mathbf{MV_{x}};x_{i},y_{i})$
      $(mv_{y})_{i}=\mathrm{P2G}(\mathbb{I},\mathbf{MV_{y}};x_{i},y_{i})$
@@
      Particles strain rate tensor (velocity gradient) is computed as the G2PD (19) of the ratio between grid nodes momentum and mass, and then particles properties are updated.
      $\forall k=0,\ldots N_{p}-1$
      $\nabla(V_{x})_{k}=\mathrm{G2PD}(\theta,\mathbf{mv_{x}},\mathbf{m};X_{k},Y_{k})$
      $\nabla(V_{y})_{k}=\mathrm{G2PD}(\theta,\mathbf{mv_{y}},\mathbf{m};X_{k},Y_{k})$
      $E_{k}^{\prime}=E_{k}+\Delta t\dfrac{-(P_{k}+Q_{k})}{\rho_{k}}\nabla\cdot\mathbf{V}_{k}$
      $\rho_{k}^{\prime}=\frac{\rho_{k}}{1+\Delta t\nabla\cdot\mathbf{V}_{k}}$
      $P_{k}^{\prime}=(\gamma-1)\rho_{k}^{\prime}E_{k}^{\prime}$
@@
      Particles position are updated using $\mathbf{\overline{V}}$:
      $\forall k=0,\ldots N_{p}-1$
      $(X_{k})^{\prime}=X_{k}+\Delta t(\overline{V_{x}})_{k}$
      $(Y_{k})^{\prime}=Y_{k}+\Delta t(\overline{V_{y}})_{k}$
      Thus, the time advancing scheme is the same of [13]. However, particles moving is postponed w.r.t. the standard algorithm [13], so that, in all the above steps, basis functions and their gradient can be computed on the fly, without the need to store the value they had at former particles position.
@@
It is worth noting that, even though the sum in the analytical expression of the Grid-to-Particle kernels is performed on the grid index, in the implementation also G2P and G2PD kernels are performed on particles index, and, internally, they rely on the particle-to-grid map, as shown in the snippet in Fig. 1. Indeed, the implementation has been found to be more efficient when the Lagrangian character of the method is taken into account. In particular, in the parallelisation scheme one thread is requested for each particle.
Algorithm 1 summarises the for loop procedure, with a simplified notation.
@@
$\qquad m=\mathrm{P2G}(M),\ \mathbf{mv}=\mathrm{P2G}(M\mathbf{V}),\ \mathbf{f}=\mathrm{P2GD}((P+Q)\tfrac{M}{\rho})$
4:     Boundary conditions enforcement on $\mathbf{mv},\,\mathbf{f}$ obstacle treatment as in (21), (22)
@@
$\qquad(\mathbf{mv})^{\prime}=\mathbf{mv}+\Delta t\mathbf{f}$
@@
$\qquad\overline{\mathbf{V}}=\mathrm{G2P}(\tfrac{\mathbf{mv}}{m}),\ {\mathbf{A}}=\mathrm{G2P}(\tfrac{\mathbf{f}}{m})$
@@
$\qquad\mathbf{V}^{\prime}=\mathbf{V}+\Delta t\cdot\mathbf{A}$